OpenStudy (eyb0ss8):

Circle A has a diameter of 8 inches, a circumference of 25.12 inches, and an area of 50.24 square inches. The diameter of circle B is 3 inches, the circumference is 9.42 inches, and the area is 7.065 square inches. Part A: Using the formula for circumference, solve for the value of pi for each circle. (4 points) Part B: Use the formula for area and solve for the value of pi for each circle. (4 points) Part C: What observation can you make about the value of pi for circles A and B? (2 points) help?

OpenStudy (phi):

it says "Using the formula for circumference, " Do you know the formula for find the circumference of a circle ?

OpenStudy (eyb0ss8):

yeah

OpenStudy (phi):

what is it ?

OpenStudy (eyb0ss8):

c=3.14r squared

OpenStudy (phi):

almost. it's not 3.14

OpenStudy (eyb0ss8):

ooo

OpenStudy (eyb0ss8):

hold on

OpenStudy (phi):

also, that is the formula for the area. you want the formula for the circumference.

OpenStudy (eyb0ss8):

2(3.14)r

OpenStudy (phi):

It's a famous formula. \[ C= 2 \pi r \] we use the Greek letter \( \pi\) because if we tried to write out the actual number, it goes on forever.

OpenStudy (phi):

in other words, 3.14 is just "roughly" the actual value of pi now they say Circle A has a diameter of 8 inches what is the radius of circle A? any idea ?

OpenStudy (phi):

ok, except use pi instead of 3.14 pi * 8 = 25.12 now "solve for pi" by dividing both sides by 8 pi= 24.12/8 what do you get ?

OpenStudy (phi):

you get pi= 24.12/8 = 3.24 and if you do the same for circle B: pi * 3= 9.42 pi = 9.42/3 = 3.14 so you get the same number for both circles. That is part 1

OpenStudy (phi):

*part A for part B, write down the formula for the area of a circle (and use pi)

OpenStudy (phi):

A= pi r^2 now put in the area for circle A 50.24 and the radius (it is 4) 50.24= pi * 4*4 or 50.24= pi * 16 solve for pi (divide both sides by 16)

OpenStudy (phi):

can you do the same for the other circle (using its area) ?

OpenStudy (phi):

Part B: Use the formula for area and solve for the value of pi for each circle. (4 points)

OpenStudy (phi):

this question is different from what you did. they want you to "solve for pi" using the info that the area of circle B is 7.065 square inches. also, they give you B's diameter as 3. you need B's radius

OpenStudy (phi):

you start with \[ A = \pi r^2\] put in 7.065 for the area (replace A with that number) also, put in the radius then "solve for pi"

OpenStudy (phi):

the radius is 1/2 of 3 or 3/2 or 1.5

OpenStudy (phi):

in other words, they told you the diameter is 3, so you know the radius is half of that

OpenStudy (phi):

put in 7.065 for the area (replace A with that number) also, put in the radius then "solve for pi"

OpenStudy (phi):

we are doing part B Part B: Use the formula for area and solve for the value of pi for each circle. for circle A you did A= pi r^2 now put in the area for circle A 50.24 and the radius (it is 4) 50.24= pi * 4*4 or 50.24= pi * 16 solve for pi (divide both sides by 16) pi = 50.24/16 = 3.14 we found pi is 3.14 for circle A now we are doing circle B A= pi r^2 put in the area for circle B 7.065 and the radius (it is 1.5)

OpenStudy (phi):

can you show your work ?

OpenStudy (phi):

ok, but the formula is 7.065 = pi * (1.5)^2 1.5^2 means 1.5*1.5

OpenStudy (eyb0ss8):

so pi*1.5*1.5=7.065 right?